10. Find the eigen values and eigen vectors of the given matrix 11 30 ( 36)
Answer #1

-- 10 12 1 Sopi 1•1: 30] . det A = -6-30--36 Characteristic is 2²+57 36= or, a2 +99- 49 -36 -0 er, a (+9)- 4(+9) = 0 or (2-4)